OpenStudy (anonymous):

Suppose that $1200 is initially invested in an account at a fixed interest rate, compounded continuously. Suppose also that, after five years, the amount of money in the account is $1360 . Find the interest rate per year? Write your answer as a percentage.

OpenStudy (anonymous):

Formula for continuously compounded interest: Amount = Pe^(rt) Plug everything in... 1360 = 1200e^(r*5) Divide by 1200. 1.13333333 = (e^5r) Natural log of both sides. 5r = 0.12516314 Divide by 5. r = 0.025032628 Number would look pretty clean at 2.5% so I'm plug it in and see if that fixes rounding errors. Amount = Pe^(rt) 1360 = 1200e^(5*0.025) 1360 approx = 1 359.77814 It didn't. So, the rate is approximately 2.5032628%

OpenStudy (anonymous):

thank u all

OpenStudy (anonymous):

@campbell_st That isn't compounded continuously.